RawlsK29:

Which of these is a simplified form of the equation 8y + 4 = 6 + 2y + 1y? 5y = 2 5y = 10 11y = 2 11y = 10

Shadow:

\[8y + 4 = 6 + 2y + 1y\] Combine like terms \[8y + 4 = 3y + 6\] Subtract 3y from both sides Subract 4 from both sides \[5y = 2\]

RawlsK29:

Thank You!!😃

Shadow:

No problem